A man is 8 times older than his son. After 5 years, he will only be 3 times older than his son. How old are they now?
Answer by stanbon(75887)

A man is 8 times older than his son. After 5 years, he will only be 3 times older than his son. How old are they now?
----
Equations:
f = 8s
(f+5)=3(s+5)
----
Substitute for "f" and solve for "s":
(8s+5) = 3s+15
5s = 10
son = 2 yrs (Son's age now)
---
Solve for "f":
f = 8s
father = 16 yrs (Father's age now
====================================
Cheers,
Stan H.